# SPECIAL RELATIVITY

### v9

@@ 9.2 Equivalence Principle and the Gravitational Physics
We should make one further remark about the equivalence principle (1). The widely held opinion that such a principle is the basis of the general relativity is not entirely correct. The gravitational field is not simply a force that is universal and has the same action on all bodies. According to Einstein’s idea, the gravitational field is associated with the structure of space-time. The first thing we must understand is the Riemann tensor, since it is this tensor that is the gravitational field: if it vanishes (and only in this case) a field does not exist. Actually this argument is already fully sufficient to show that the equivalence principle in the usual formulation does not hold. The equivalence principle is valid for mechanical processes and does not hold for electrodynamics processes. This means that by means of measurements within the a system it is possible to establish whether the system is inertial or falls freely in a homogeneous gravitational field. In a homogeneous gravity field, a free charge is uniformly accelerated with respect to an inertial frame, and therefore the charge will radiate, and radiation will occur in both strong and weak fields. In the uniformly accelerated frame of reference, the free charge is not accelerated with respect to the inertial frame of reference. But in then follows in accordance with electrodynamics that it can never radiate, and therefore the equivalence principle is indeed violated .
@@
With the special and general relativity we can discuss the case of earth-based experiments where pseudo-gravity and the real gravity are inextricably mixed. With this understanding, we next need to describe the gravitational fields near the sun due to the sun mass itself. Let us consider the simplest instance of a transformation from an inertial frame, in which the space-time is Minkowskian, to the earth-based frame. Newton gravity is known to be valid when gravitation fields are to weak to produce velocities near the speed of light. In other words, the gravitational potential energy of a particle is much less than its rest-mass energy. In such situation, general relativity must make the same predictions as Newton gravity. For weak gravitational field the ordinary Newtonian potential $\phi$ completely determines the metric, which has the form $ds^{2}=(1+2\phi/c^{2})c^{2}dt^{2}-(1-2\phi/c^{2})(dx^{2}+dy^{2}+dz^{2})$. The sign of $\phi$ is chosen negative, so that far from the source of Mass $M$, we have $\phi=-GM/r$. Now the condition above that the field be weak means that $|m\phi|<<mc^{2}$, so that $|\phi/c^{2}|<<1$. This metric is really only correct to first order in $\phi/c^{2}$. We say that space-time becomes flat asymptotically at large distances from the source. The Langevin metric for rotating frame could be used for rotating (around the sun) earth: $ds^{2}=(1-v^{2}/c^{2})d(ct)^{2}-2(\vec{v}\cdot d\vec{r})dt-(d\vec{r})^{2}$. Where $\vec{v}$ is the velocity of the earth, and $(d\vec{r})^{2}$ is the square of the coordinate distance. The absence of equivalence of the equations in an accelerated reference system and in an inertial frame with a static gravitational field is obvious, since in the first case the non-diagonal metric coefficients will be functions of the time, whereas in the second case the metric exhibits diagonal properties and does not depend on the time. Therefore, in the first case the charge at rest will radiate, whereas in the second case the charge at rest will not radiate. It is for this reason that one cannot deduce from the observations the validity of the equivalence principle as a physical principle. The relation between gravity and inertia is a subject that needs to be understood more deeply. Now let us return to observations of an earth-based observer. One additional expression for the invariant interval is needed: $ds^{2}=(1+2\phi/c^{2}-v^{2}/c^{2})d(ct)^{2}-2(\vec{v}\cdot d\vec{r})dt-(1-2\phi/c^{2})(d\vec{r})^{2}$. To the order of calculations, this result is a simple superposition of the Langevin metric with the metric in the sun centered inertial frame for the weak external gravitational field of the sun. Both orbit speed and gravitational field dilate time and contract space. The differences are small, however, and the Langevin metric represents an extremely good approximation for the metric in the rotating earth-based reference system.
@@ 9.3 Time Dilation
It is generally believed that the special relativity is a reciprocal theory. An excellent illustration of the absence of reciprocity is the time dilation effect. Consider the course of time in two inertial reference frames, one of which, $S$, will be considered to be at rest, while another one, $S_{n}$, will accelerate with respect to the first one up to velocity $v$ along the $x$- axis. As we already discussed, the transformation from the inertial frame to the accelerated frame apparently first used by Langevin. The mathematical idea that are used in the literature of time dilation effect is the smooth tailoring of the metric tensor. The transformation to the accelerated frame, with coordinates denoted by ”n”, has the Galilean form, $t_{n}=t$, $x_{n}=x-vt$. After the Galilean transformation, the metric of non-inertial frame goes to into the Langevin metric of the accelerated inertial frame $S_{n}$ in a continuous manner (see Section 9.6 for more detail).
@@
Suppose we have two identical physical clocks at one and the same point of the inertial reference system $S$. Consider their readings to coincide at the initial moment $t=0$. Let first of these clocks always be at rest in the frame $S$. At moment $t=0$ the second clock accelerates and starts to move with a constant velocity $v$ along the $x$ axis. Now consider the reading of the clocks in the accelerated reference system, where the second clock is always at rest. Frames are not equivalent, since frame $S_{n}$ accelerated with respect to the fixed stars. In the accelerated frame the first clock moves with velocity $dx_{n}/dt_{n}=-v$. Taking into account the Langevin metric Eq.(15), $ds^{2}=c^{2}(1-v^{2}/c^{2})dt_{n}^{2}-2vdx_{n}dt_{n}-dx_{n}^{2}$, we obtain $d\tau=ds/c=dt_{n}$, i.e. the time, shown by the first physical clock will coincide with the time $t=t_{n}$ of coordinate clock. Since the second physical clock is at rest, its reading of its proper time is $d\tau_{n}=\sqrt{1-v^{2}/c^{2}}dt_{n}$. This demonstrates the time dilation of a physical clock at rest in the accelerated frame as compared to physical clocks in the inertial frame. We have found that the slowly down of the accelerated physical clock does not depend on the choice of reference system, in which this effect is measured as it must be.
@@
Now we go on to consider another situation, a very practical one. Suppose that we split the accelerated frame. Now frames $S_{n}^{(1)}$ and $S_{n}^{(2)}$ accelerate with respect to the inertial frame $S$, up to the same speed $v$ but move in opposite directions. The problem is complete symmetrical in the inertial frame $S$ and no difference in the time dilation of a physical clock at rest in the accelerated frames $S_{n}^{(1)}$ and $S_{n}^{(2)}$ (as compared to physical clock in the inertial frame $S$) can exist. The following simple analysis confirms this conclusion. Suppose we have two physical clocks. Let first of these clocks be at rest in the frame $S_{n}^{(1)}$. Now consider the reading of the second clock in the accelerated system $S_{n}^{(2)}$ where the second clock is at rest. In the accelerated frame $S_{n}^{(1)}$ second clock moves with velocity $dx_{n}/dt_{n}=-2v$. Taking into account the Langevin metric, $ds^{2}=c^{2}(1-v^{2}/c^{2})dt_{n}^{2}-2vdx_{n}dt_{n}-dx_{n}^{2}$, we obtain $d\tau_{2}=ds/c=\sqrt{1-v^{2}/c^{2}}dt_{n}$, i.e. the time, shown by the first physical clock $d\tau_{1}$ coincide with the physical time $d\tau_{2}$ of the second physical clock. The situation seems paradoxical. Now the accelerated frames $S_{n}^{(1)}$ and $S_{n}^{(2)}$ move relative to each other with a non zero speed. We see an interesting thing: the presence of some non zero relative velocity cannot be a cause for the effect of time dilation.
@@
Above we discussed that one can describe the non-inertial frames using the language of general relativity. Time dilation effect associated with the transformation from inertial frame to the accelerated frame may be regarded as a result of the action of certain force. A resolution of the symmetry paradox identifies a pseudo-gravitational field within the systems $S_{n}^{(1)}$ and $S_{n}^{(2)}$ as the agency of time dilation effect. Proper time clocks having the same accelerational histories really do run at same rates and yield identical measurement results when at rest in different inertial systems.
@@
Now we go on to consider another situation. Suppose that frame $S_{n}$ accelerates with respect to the inertial frame $S$ up to the speed $v$. Now frames $S_{n}^{(1)}$ and $S_{n}^{(2)}$ accelerate with respect to the frame $S_{n}$ up to the same speed $v$ but move in opposite direction. The problem is not symmetrical in the frame $S_{n}$ and there is a difference in the time dilation in the accelerated frames $S_{n}^{(1)}$ and $S_{n}^{(2)}$. In the frame $S_{n}$ the frame $S_{n}^{(1)}$ moves with velocity $-v$ and we obtain proper time $d\tau_{n}^{(1)}=dt=dt_{n}$. In the frame $S_{n}$ second frame moves with velocity $v$. Taking into account the Langevin metric, $ds^{2}=c^{2}(1-v^{2}/c^{2})dt_{n}^{2}-2vdx_{n}dt_{n}-dx_{n}^{2}$, we obtain $d\tau_{n}^{(2)}=\sqrt{1-4v^{2}/c^{2}}dt_{n}$. Thus we observe time dilation of the second frame with respect to the inertial frame $S$. We have made this analysis from the point of view of an observer at rest in the frame $S_{n}$: we would like to know how it would look to the observer who is at rest in the inertial frame. If we analyze time dilation in the frame $S$ we find that frame $S_{n}^{(2)}$ moves with velocity $dx/dt=2v$. Taking into account the Minkowski metric we obtain $d\tau_{n}^{(2)}=\sqrt{1-4v^{2}/c^{2}}dt$ as it must be. In our calculations we find that mathematics for relating proper times is extremely simple. That is not too surprising. We used Galilean boosts and the Galilean addition of velocities. We shall discuss the time dilation effect for moving reference (atomic) clock further in the Chapter 11.
@@
9.4.1 Introductory remarks
In order to understand the slowing of the clock in an accelerated system, we have to watch the machinery of the clock and see what happens when it is moving. We shall try to understand the effect in a very simple case. A clock which we shall call ”light-clock” is a rod (meter stick) with a mirror at each end. The tick of time is the reflection of light in the mirror. Let us now consider the parallel and orthogonal to motion orientations of the light clock. The equivalence of all clocks underlines the theory of relativity and we shall show that irrespective of the orientation of the light path with respect to the motion of the light-clock, the clock scores the same time. Two light-clocks are shown in Fig. 46
@@ 9.4.1 Introductory remarks
Now we have to watch the machinery of the physical clock and see what happens to the accelerated light-clock. We describe the clock based on the observations made by an observer in the same accelerated frame as the clock. In a first step, we set up the light clock such that light travels between mirrors parallel to the direction of fixed stars motion. In the case of Langevin metric, the speed of light radiated by the accelerated emitter cannot be equal $c$ anymore. If $ds$ is the infinitesimal displacement along the world line of a beam of light, then $ds^{2}=0$ and using Eq.(15) we obtain $c^{2}=(dx_{n}/dt_{n}+v)^{2}$. This means that in the accelerated reference system of coordinates $(ct_{n},x_{n})$ the velocity of light radiated by the accelerated emitter parallel to the $x_{n}$-axis, is $dx_{n}/dt_{n}=c-v$ in the positive direction, and $dx_{n}/dt_{n}=-c-v$ in the negative direction. Designating the rod (meter stick) length, unknown so far, by $l_{n}$, we obtain the time in which light gets from the left mirror to the right mirror, $t_{1}=l_{n}/(c-v)$, and the time in which light gets from the right mirror to the left mirror, $t_{2}=l_{n}/(c+v)$. Therefore, the time interval between the sending and reception of the light signal is $\Delta t=t_{1}+t_{2}=2l_{n}/[c(1-v^{2}/c^{2})]$.
One must first determine the length of physical rods. For the measurement of distances between fixed points on the accelerated frame we shall use physical measuring rods of the same kind as those used in the inertial frame, but now at rest relative to the accelerated frame. In connection with the process of measuring distances in accelerated frames a problem arises which did not occur in inertial frames. For according to the special relativity, no rigid bodies can exist, since they would provide a means of transmitting signals with velocity greater than velocity of light.
@@
With regard to spatial distances, the interpretive principle is that $\sqrt{-ds^{2}}$ gives the length of an infinitesimal rod whose endpoints are simultaneous according to standard simultaneity in the rod’s rest frame. Taking into account the Langevin metric Eq.(15), we obtain at $dt_{n}=0$ that $dx_{n}=dx$, i.e. the length of physical rod in the inertial frame will coincide with the coordinate distance in the accelerated frame. When we apply this rule to physical rods that are at rest in the accelerated frame, we encounter the complication that $dt_{n}=0$ does not automatically correspond to standard simultaneity in the accelerated frame. The length will surely be different for a physical rod at rest in the accelerated frame compared to a physical rod at rest in the inertial frame.
@@
The spatial geometry in the accelerated frame is characterized by the differential spatial line element on it. It should be note that spatial geometry has a conventional character. However, the proper spatial line element $dl_{n}$ is of special significance, since it defines a coordinate-independent spatial geometry in the accelerated frame. The relation $dl_{n}^{2}=(-g_{\alpha\beta}+g_{0\alpha}g_{0\beta}/g_{00})dx_{\alpha}dx_{\beta}$ gives the connection between the 3-dimensional spatial line element and the metric of the four-dimensional space-time. It can be shown that the element $dl_{n}$ is invariant under a transformation connecting two different coordinate systems inside the same frame of reference. We note that the spatial line element $dl$ obtained by putting $dt_{n}=0$ in Eq.(15), $dl^{2}=dx_{n}^{2}$, is differ from $dl_{n}$.
@@
With the help of Eq.(15), we find $dl_{n}^{2}=dx_{n}^{2}/(1-v^{2}/c^{2})$, which determines the spatial geometry in the accelerated frame. This equation is asymmetric with respect to the lengths $dx_{n}=dl$ and $dl_{n}$, since it relates the physical length $dl$ of the rod in the inertial frame $S$ to the physical length $dl_{n}$ in the accelerated frame $S_{n}$. Our analysis based on using standard measuring rods in an accelerated frame to measure it geometrical properties shows that coordinate distance $dx_{n}$ has physical length $dx_{n}/\sqrt{1-v^{2}/c^{2}}$. With reference to the accelerated frame it is interpreted as an effect of acceleration with respect to the fixed stars, i.e. the geometry of space is non-Euclidean in an accelerated frame.
@@
This shows that the physical meter stick on which the mirrors are mounted undergoes compress on in the direction of fixed stars motion. Thus we have ascertained the asymmetry of meter stick’s length directly from the metric. Therefore, the time interval between the sending and reception of the light signal in the parallel light-clock in the accelerated frame is $\Delta t_{n}=(2l/c)/\sqrt{1-v^{2}/c^{2}}$. Because $t_{n}=t$ and $\Delta t=2l/c$ has the physical meaning of the time indicated by a physical clock at rest in the inertial frame, this implies that physical clocks at rest in the accelerated frame are slow compared to physical clocks in the original inertial frame.
9.4.3 Orthogonal Light-Clock. Langevin Metric
@@
Let us consider a light-clock accelerated up to velocity $v$ transverse to the direction of optical pulse. In other words, we set up light-clock such that the effect of a meter stick contraction is not present. We describe the orthogonal light-clock based on the observations made by an observer in the same accelerated frame as the light-clock. If $ds$ is the infinitesimal displacement along the world line of a beam of light, then $ds^{2}=0$ and using Eq.(15) we obtain $(dz_{n}/dt_{n})^{2}=c^{2}(1-v^{2}/c^{2})$. Therefore, the time interval between the sending and reception of the light signal is $\Delta t_{n}=2l/[c\sqrt{1-v^{2}/c^{2}}]$. Because $t_{n}=t$ and $\Delta t=2l/c$ has the physical meaning of the time indicated by a physical clock at rest in the inertial frame, this implies that physical clocks at rest in the accelerated frame are slow compared to physical clocks in the original inertial frame. The theory of relativity shows us that irrespective of the orientation of the light path with respect to the motion of the light-clock, the clock scores the same time.

### v10

@@ 12.2 Equivalence Principle and the Gravitational Physics
We should make one further remark about the equivalence principle  [^52] For a general discussion of the equivalence principle we suggest reading the paper [50].. The widely held opinion that such a principle is the basis of general relativity is not entirely correct. The gravitational field is not simply a force that is universal and has the same action on all bodies. According to Einstein’s idea, the gravitational field is associated with the structure of space-time. The first thing we must understand is the Riemann tensor, since it is this tensor that is the gravitational field: if it vanishes (and only in this case) a field does not exist. Actually, this argument is already fully sufficient to show that the equivalence principle in the usual formulation does not hold. The equivalence principle is valid for mechanical processes and does not hold for electrodynamics processes. This means that by means of measurements within the system it is possible to establish whether the system is inertial or falls freely in a homogeneous gravitational field. In a homogeneous gravity field, a free charge is uniformly accelerated with respect to an inertial frame, and therefore the charge will radiate, and radiation will occur in both strong and weak fields. In the uniformly accelerated frame of reference, the free charge is not accelerated with respect to the inertial frame of reference. But in accordance with electrodynamics, it can never radiate, and therefore the equivalence principle is indeed violated.
@@
With special and general relativity we can discuss the case of earth-based experiments where pseudo-gravity and real gravity are inextricably mixed. With this understanding, we next need to describe the gravitational fields near the sun due to the sun’s mass itself. Let us consider the simplest instance of a transformation from an inertial frame, in which the space-time is Minkowskian, to the earth-based frame. Newton’s gravity is known to be valid when gravitation fields are too weak to produce velocities near the speed of light. In other words, the gravitational potential energy of a particle is much less than its rest-mass energy. In such a situation, general relativity must make the same predictions as Newton gravity. For weak gravitational field, the ordinary Newtonian potential $\phi$ completely determines the metric, which has the form $ds^{2}=(1+2\phi/c^{2})c^{2}dt^{2}-(1-2\phi/c^{2})(dx^{2}+dy^{2}+dz^{2})$. The sign of $\phi$ is chosen negatively so that far from the source of Mass $M$, we have $\phi=-GM/r$. Now the condition above that the field be weak means that $|m\phi|<<mc^{2}$, so that $|\phi/c^{2}|<<1$. This metric is only correct to the first order in $\phi/c^{2}$. We say that space-time becomes flat asymptotically at large distances from the source. The Langevin metric for rotating frame could be used for rotating (around the sun) earth: $ds^{2}=(1-v^{2}/c^{2})d(ct)^{2}-2(\vec{v}\cdot d\vec{r})dt-(d\vec{r})^{2}$. Where $\vec{v}$ is the velocity of the earth, and $(d\vec{r})^{2}$ is the square of the coordinate distance. The absence of equivalence of the equations in an accelerated reference system and in an inertial frame with a static gravitational field is obvious since in the first case the non-diagonal metric coefficients will be functions of the time, whereas in the second case the metric exhibits diagonal properties and does not depend on the time. Therefore, in the first case, the charge at rest will radiate, whereas in the second case, the charge at rest will not radiate. It is for this reason that one cannot deduce from the observations the validity of the equivalence principle as a physical principle. Now let us return to observations of an earth-based observer. One additional expression for the invariant interval is needed: $ds^{2}=(1+2\phi/c^{2}-v^{2}/c^{2})d(ct)^{2}-2(\vec{v}\cdot d\vec{r})dt-(1-2\phi/c^{2})(d\vec{r})^{2}$. To the order of calculations, this result is a simple superposition of the Langevin metric with the metric in the sun-centered inertial frame for the weak external gravitational field of the sun. Both orbit speed and gravitational field dilate time and contract space. The differences are small, however, and the Langevin metric represents an extremely good approximation for the metric in the rotating earth-based reference system.
@@ 12.3 Time Dilation
It is generally believed that the special relativity is a reciprocal theory. An excellent illustration of the absence of reciprocity is the time dilation effect. Consider the course of time in two inertial reference frames, one of which, $S$, will be considered to be at rest with respect to the fixed stars, while another one, $S_{n}$, will accelerate with respect to the first one up to velocity $v$ along the $x$-axis. As we already discussed, the transformation from the inertial frame to the accelerated frame was apparently first used by Langevin. The mathematical idea used in the literature on the time dilation effect is the smooth tailoring of the metric tensor. The transformation to the accelerated frame, with coordinates denoted by ”n”, has the Galilean form, $t_{n}=t$, $x_{n}=x-vt$. After the Galilean transformation, the metric of the non-inertial frame goes to the Langevin metric of the accelerated inertial frame $S_{n}$ in a continuous manner (see Chapter 5 for more detail).
@@
Suppose we have two identical physical clocks at the same point of the inertial reference system $S$. Consider their readings to coincide at the initial moment $t=0$. Let first of these clocks always be at rest in the frame $S$. At moment $t=0$ the second clock accelerates and starts to move with a constant velocity $v$ along the $x$ axis. Now consider the reading of the clocks in the accelerated reference system, where the second clock is always at rest. Frames are not equivalent, since frame $S_{n}$ accelerated with respect to the fixed stars. In the accelerated frame the first clock moves with velocity $dx_{n}/dt_{n}=-v$. Taking into account the Langevin metric Eq.(15), $ds^{2}=c^{2}(1-v^{2}/c^{2})dt_{n}^{2}-2vdx_{n}dt_{n}-dx_{n}^{2}$, we obtain $d\tau=ds/c=dt_{n}$, i.e. the time, shown by the first physical clock will coincide with the time $t=t_{n}$ of coordinate clock. Since the second physical clock is at rest, the reading of its proper time is $d\tau_{n}=\sqrt{1-v^{2}/c^{2}}dt_{n}$. This demonstrates the time dilation of a physical clock at rest in the accelerated frame as compared to physical clocks in the inertial frame. We have found that the slowdown of the accelerated physical clock does not depend on the choice of reference system, in which this effect is measured.
@@
Now we go on to consider another situation, a very practical one. Suppose that we split the accelerated frame. Now frames $S_{n}^{(1)}$ and $S_{n}^{(2)}$ accelerate with respect to the inertial frame $S$, up to the same speed $v$ but move in opposite directions. The problem is completely symmetrical in the inertial frame $S$ and no difference in the time dilation of a physical clock at rest in the accelerated frames $S_{n}^{(1)}$ and $S_{n}^{(2)}$ (as compared to physical clock in the inertial frame $S$) can exist. The following simple analysis confirms this conclusion. Suppose we have two physical clocks. Let first of these clocks be at rest in the frame $S_{n}^{(1)}$. Now consider the reading of the second clock in the accelerated system $S_{n}^{(2)}$ where the second clock is at rest. In the accelerated frame $S_{n}^{(1)}$ second clock moves with velocity $dx_{n}/dt_{n}=-2v$. Taking into account the Langevin metric, $ds^{2}=c^{2}(1-v^{2}/c^{2})dt_{n}^{2}-2vdx_{n}dt_{n}-dx_{n}^{2}$, we obtain $d\tau_{2}=ds/c=\sqrt{1-v^{2}/c^{2}}dt_{n}$, i.e. the time, shown by the first physical clock $d\tau_{1}$ coincide with the physical time $d\tau_{2}$ of the second physical clock. The situation seems paradoxical. Now the accelerated frames $S_{n}^{(1)}$ and $S_{n}^{(2)}$ move relative to each other with a non-zero speed. We see an interesting thing: the presence of some non-zero relative velocity cannot be a cause for the effect of time dilation.
@@
Above we discussed that one can describe the non-inertial frames using the language of general relativity. The time dilation effect associated with the transformation from the inertial frame to the accelerated frame may be regarded as a result of the action of a certain force. A resolution of the symmetry paradox identifies a pseudo-gravitational field within the systems $S_{n}^{(1)}$ and $S_{n}^{(2)}$ as the agency of the time dilation effect. Proper time clocks having the same accelerational histories do run at the same rates and yield identical measurement results when at rest in different inertial systems.
@@
Now we go on to consider another situation. Suppose that frame $S_{n}$ accelerates with respect to the inertial frame $S$ up to the speed $v$. Now frames $S_{n}^{(1)}$ and $S_{n}^{(2)}$ accelerate with respect to the frame $S_{n}$ up to the same speed $v$ but move in opposite direction. The problem is not symmetrical in the frame $S_{n}$ and there is a difference in the time dilation in the accelerated frames $S_{n}^{(1)}$ and $S_{n}^{(2)}$. In the frame $S_{n}$ the frame $S_{n}^{(1)}$ moves with velocity $-v$ and we obtain proper time $d\tau_{n}^{(1)}=dt=dt_{n}$. In the frame $S_{n}$ second frame moves with velocity $v$. Taking into account the Langevin metric, $ds^{2}=c^{2}(1-v^{2}/c^{2})dt_{n}^{2}-2vdx_{n}dt_{n}-dx_{n}^{2}$, we obtain $d\tau_{n}^{(2)}=\sqrt{1-4v^{2}/c^{2}}dt_{n}$. Thus we observe time dilation of the second frame with respect to the inertial frame $S$. We have made this analysis from the point of view of an observer at rest in the frame $S_{n}$: we would like to know how it would look to the observer who is at rest in the inertial frame. If we analyze time dilation in the frame $S$ we find that frame $S_{n}^{(2)}$ moves with velocity $dx/dt=2v$. Taking into account the Minkowski metric we obtain $d\tau_{n}^{(2)}=\sqrt{1-4v^{2}/c^{2}}dt$ as it must be. In our calculations, we find that mathematics for relating proper times is extremely simple. That is not too surprising. We used Galilean boosts and the Galilean addition of velocities.
@@
To understand the slowing of the clock in an accelerated system, we have to watch the machinery of the clock and see what happens when it is moving. We shall try to understand the effect in a very simple case. A clock which we shall call ”light clock” is a rod (meter stick) with a mirror at each end. The tick of time is the reflection of light in the mirror. Let us now consider the parallel and orthogonal motion orientations of the light clock. The equivalence of all clocks underlines the theory of relativity and we shall show that irrespective of the orientation of the light path with respect to the motion of the light-clock, the clock scores the same time. Two light clocks are shown in Fig. 47
@@
Now we have to watch the machinery of the physical clock and see what happens to the accelerated light clock. We describe the clock based on the observations made by an observer in the same accelerated frame as the clock. In the first step, we set up the light clock such that light travels between mirrors parallel to the direction of fixed stars motion. In the case of the Langevin metric, the speed of light radiated by the accelerated emitter cannot be equal to $c$ anymore. If $ds$ is the infinitesimal displacement along the world line of a beam of light, then $ds^{2}=0$, and using Eq.(15) we obtain $c^{2}=(dx_{n}/dt_{n}+v)^{2}$. This means that in the accelerated reference system of coordinates $(ct_{n},x_{n})$ the velocity of light radiated by the accelerated emitter parallel to the $x_{n}$-axis, is $dx_{n}/dt_{n}=c-v$ in the positive direction, and $dx_{n}/dt_{n}=-c-v$ in the negative direction. Designating the rod (meter stick) length, unknown so far, by $l_{n}$, we obtain the time in which light gets from the left mirror to the right mirror, $t_{1}=l_{n}/(c-v)$, and the time in which light gets from the right mirror to the left mirror, $t_{2}=l_{n}/(c+v)$. Therefore, the time interval between the sending and reception of the light signal is $\Delta t=t_{1}+t_{2}=2l_{n}/[c(1-v^{2}/c^{2})]$.
@@
One must first determine the length of the physical rods. For the measurement of distances between fixed points on the accelerated frame we shall use physical measuring rods of the same kind as those used in the inertial frame, but now at rest relative to the accelerated frame. With regard to spatial distances, the interpretive principle is that $\sqrt{-ds^{2}}$ gives the length of an infinitesimal rod whose endpoints are simultaneous according to standard simultaneity in the rod’s rest frame. Taking into account the Langevin metric Eq.(15), we obtain at $dt_{n}=0$ that $dx_{n}=dx$, i.e. the length of the physical rod in the inertial frame will coincide with the coordinate distance in the accelerated frame. When we apply this rule to physical rods that are at rest in the accelerated frame, we encounter the complication that $dt_{n}=0$ does not automatically correspond to standard simultaneity in the accelerated frame. The length will surely be different for a physical rod at rest in the accelerated frame compared to a physical rod at rest in the inertial frame.
@@
The spatial geometry in the accelerated frame is characterized by the differential spatial line element on it. It should be noted that spatial geometry has a conventional character. However, the proper spatial line element $dl_{n}$ is of special significance, since it defines a coordinate-independent spatial geometry in the accelerated frame. The relation $dl_{n}^{2}=(-g_{\alpha\beta}+g_{0\alpha}g_{0\beta}/g_{00})dx_{\alpha}dx_{\beta}$ gives the connection between the 3-dimensional spatial line element and the metric of the four-dimensional space-time. It can be shown that the element $dl_{n}$ is invariant under a transformation connecting two different coordinate systems inside the same frame of reference. We note that the spatial line element $dl$ obtained by putting $dt_{n}=0$ in Eq.(15), $dl^{2}=dx_{n}^{2}$, is differ from $dl_{n}$.
@@
With the help of Eq.(15), we find $dl_{n}^{2}=dx_{n}^{2}/(1-v^{2}/c^{2})$, which determines the spatial geometry in the accelerated frame. This equation is asymmetric with respect to the lengths $dx_{n}=dl$ and $dl_{n}$ since it relates the physical length $dl$ of the rod in the inertial frame $S$ to the physical length $dl_{n}$ in the accelerated frame $S_{n}$. Our analysis based on using standard measuring rods in an accelerated frame to measure its geometrical properties shows that coordinate distance $dx_{n}$ has physical length $dx_{n}/\sqrt{1-v^{2}/c^{2}}$. It is interpreted as an effect of acceleration with respect to the fixed stars, i.e. the geometry of space is non-Euclidean in an accelerated frame.
@@
This shows that the physical meter stick on which the mirrors are mounted undergoes compression in the direction of fixed stars motion. Thus we have ascertained the asymmetry of the meter stick’s length directly from the metric. Therefore, the time interval between the sending and reception of the light signal in the parallel light clock in the accelerated frame is $\Delta t_{n}=(2l/c)/\sqrt{1-v^{2}/c^{2}}$. Because $t_{n}=t$ and $\Delta t=2l/c$ have the physical meaning of the time indicated by a physical clock at rest in the inertial frame, this implies that physical clocks at rest in the accelerated frame are slow compared to physical clocks in the original inertial frame.
@@
Let us consider a light-clock accelerated up to velocity $v$ transverse to the direction of the optical pulse. In other words, we set up light clock such that the effect of a meter stick contraction is not present. We describe the orthogonal light clock based on the observations made by an observer in the same accelerated frame as the light clock. If $ds$ is the infinitesimal displacement along the world line of a beam of light, then $ds^{2}=0$, and using Eq.(15) we obtain $(dz_{n}/dt_{n})^{2}=c^{2}(1-v^{2}/c^{2})$. Therefore, the time interval between the sending and reception of the light signal is $\Delta t_{n}=2l/[c\sqrt{1-v^{2}/c^{2}}]$. Because $t_{n}=t$ and $\Delta t=2l/c$ have the physical meaning of the time indicated by a physical clock at rest in the inertial frame, this implies that physical clocks at rest in the accelerated frame are slow compared to physical clocks in the original inertial frame.